OpenStudy (anonymous):

PLEASE HELP! find a 5th degree polynomial function f(x) with integer coefficients whose zeros are 1±2sqrt3, -2 and 0 where 0 is a zero of multiplicity 2. write the polynomial in the form asubnx^2+asub(n-1)x^(n-1)+...asub1x+a0 i understand that there should be x^2(x+2) but i'm confused with the negative signs of 1±2sqrt 3... should it be (x-1-2sqrt3) and (x-1+2sqrt3)?? or (x-1+2sqrt3) and (x+1-2sqrt3) ???

OpenStudy (anonymous):

please explain tooo!!:(

hartnn (hartnn):

ok , when the roots are \(a\pm \sqrt b\) then the factors are (x-a+\(\sqrt b\)) and (x-a-\(\sqrt b\)). does this help ?

hartnn (hartnn):

@dydlf so your first try was correct.

OpenStudy (anonymous):

thank you @hartnn !!!!!!!